OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). (explain show steps)

OpenStudy (anonymous):

3x + y = 7 y = -3x + 7 m1 = -3 Perpendicular m1*m2 = -1 -3 * m2 = -1 m2 = 1/3 Nw using point Slope Form.... (6, −1) y-y1 = m2(x-x1).....Hope nw u can Solve..

OpenStudy (anonymous):

y+1 = 1/3 ( x-6).......Just Solve.....)

OpenStudy (anonymous):

confused? y =1/3x − 3 y =1/3x + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

y =1/3x − 3

OpenStudy (anonymous):

thank you and how did you get that?